# Three-term Machin-type formulae

### v4

@@ 5. Proof of the Theorem
$[10^{7},10^{14}]$ $\geq e^{40000}$ $(10^{7},e^{40000})$ $144270$ $1.963\times 10^{12}$ $6.625\times 10^{18}$
$[5000,10^{7}]$ $\geq e^{40000}$ $(5000,e^{40000})$ $106130$ $9.900\times 10^{11}$ $4.568\times 10^{18}$
$[53,5000]$ $\geq e^{40000}$ $(53,e^{40000})$ $91595$ $7.602\times 10^{11}$ $5.926\times 10^{18}$
$\geq 5000$ $[e^{2000},e^{40000}]$ $(5000,e^{2000})$ $40000$ $1.448\times 10^{12}$ $7.586\times 10^{18}$
$[53,5000]$ $[e^{2000},e^{40000}]$ $(53,e^{2000})$ $40000$ $1.664\times 10^{11}$ $1.300\times 10^{17}$
$\geq 53$ $[m_{1}^{*},e^{2000}]$ $(53,m_{1}^{*})$ $2000$ $1.798\times 10^{12}$ $2.835\times 10^{18}$
$41$ $\geq e^{40000}$ $(41,e^{40000})$ $77848$ $2.411\times 10^{11}$ $1.608\times 10^{18}$
$37$ $\geq e^{40000}$ $(37,e^{40000})$ $71787$ $8.819\times 10^{10}$ $2.401\times 10^{17}$
$29$ $\geq e^{40000}$ $(29,e^{40000})$ $74759$ $1.461\times 10^{11}$ $6.776\times 10^{17}$
$17$ $\geq e^{40000}$ $(17,e^{40000})$ $72529$ $9.837\times 10^{10}$ $3.768\times 10^{17}$
$13$ $\geq e^{40000}$ $(13,e^{40000})$ $76708$ $1.949\times 10^{11}$ $1.544\times 10^{18}$
$5$ $\geq e^{40000}$ $(5,e^{40000})$ $76142$ $1.432\times 10^{10}$ $1.338\times 10^{18}$
@@
Thus we conclude that $x_{i}<\exp(5.306\times 10^{12})$ and $\left\lvert y_{i}\right\rvert<1.472\times 10^{19}$ in Case II.

### v5

@@ 5. Proof of the Theorem
$[10^{7},10^{14}]$ $\geq e^{40000}$ $(10^{7},e^{40000})$ $219940$ $2.992\times 10^{12}$ $1.010\times 10^{19}$
$[5000,10^{7}]$ $\geq e^{40000}$ $(5000,e^{40000})$ $161783$ $1.507\times 10^{12}$ $6.964\times 10^{18}$
$[53,5000]$ $\geq e^{40000}$ $(53,e^{40000})$ $139523$ $1.158\times 10^{12}$ $9.027\times 10^{18}$
$\geq 5000$ $[e^{2000},e^{40000}]$ $(5000,e^{2000})$ $40000$ $1.475\times 10^{12}$ $7.736\times 10^{18}$
$[53,5000]$ $[e^{2000},e^{40000}]$ $(53,e^{2000})$ $40000$ $1.664\times 10^{10}$ $1.300\times 10^{17}$
$\geq 53$ $[m_{1}^{*},e^{2000}]$ $(53,m_{1}^{*})$ $2000$ $1.830\times 10^{11}$ $2.890\times 10^{18}$
$41$ $\geq e^{40000}$ $(41,e^{40000})$ $118600$ $3.674\times 10^{11}$ $2.451\times 10^{18}$
$37$ $\geq e^{40000}$ $(37,e^{40000})$ $109518$ $1.346\times 10^{11}$ $3.663\times 10^{17}$
$29$ $\geq e^{40000}$ $(29,e^{40000})$ $113953$ $2.227\times 10^{11}$ $1.034\times 10^{18}$
$17$ $\geq e^{40000}$ $(17,e^{40000})$ $110578$ $1.501\times 10^{11}$ $5.745\times 10^{17}$
$13$ $\geq e^{40000}$ $(13,e^{40000})$ $116817$ $2.968\times 10^{11}$ $2.352\times 10^{18}$
$5$ $\geq e^{40000}$ $(5,e^{40000})$ $115906$ $2.180\times 10^{11}$ $2.038\times 10^{18}$
@@
Thus we conclude that $x_{i}<\exp(5.07\times 10^{12})$, and $\left\lvert y_{i}\right\rvert<1.316\times 10^{19}$ in Case II.